OpenStudy (anonymous):

Solve (Sqrt9x)+4=-8. Check your solution. If there is no solution, write no solution. Show your work. (Hint: Show your work when solving the equation AND when checking your solution.)

OpenStudy (anonymous):

sqrt (9x + 4) = -8 squaring both side 9x + 4= 64 9x = 60 x = 60 / 9 x = 6.666 ...... :)

OpenStudy (anonymous):

It is Sqrt'ing the 9x, then adding 4

OpenStudy (anonymous):

ok soo solve this equation, Sqrt of 9x, Then add 4, which equals -8, Then check your answer

OpenStudy (anonymous):

\[\sqrt{9x}+4=-8\] \[3\sqrt{x}=-12\] \[\sqrt{x}=-4\] \[x=16\]

OpenStudy (anonymous):

hhhhmmmm if I did it that way, squaring 9 would be 3+4=7 not -8, hhhhmmmm, idk :( sorry :(

OpenStudy (anonymous):

Its k, Eliza27 Thats good, Thank you both!!!

OpenStudy (anonymous):

That's wrong

OpenStudy (anonymous):

ur welcome for trying lolz :D

OpenStudy (anonymous):

What Akiroso?

OpenStudy (anonymous):

????

OpenStudy (anonymous):

Do you have a domine for the question?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

0.o

OpenStudy (anonymous):

H@H@ I dunno wat juss hapenn LOLZ :D :D

OpenStudy (anonymous):

Ya lol Ok im off

OpenStudy (anonymous):

thx 4 da medal :D

OpenStudy (anonymous):

thx 4 da medal :DElmoLovesU

OpenStudy (anonymous):

You can check by putting the x value back into the equation giving \[ \sqrt{9 \times16} +4=8\] just remember when you square root it you get a a positive and a negative answer